# Parameter Invariance Analysis of Moment Equations Using Dulmage-Mendelsohn Decomposition

### v1

@@ Abstract
Living organisms maintain stable functioning amid environmental fluctuations through homeostasis, a mechanism that preserves a system’s behavior despite changes in environmental conditions. To elucidate homeostasis in stochastic biochemical reactions, theoretical tools for assessing population-level invariance under parameter perturbations are crucial. In this paper, we propose a systematic method for identifying the stationary moments that remain invariant under parameter perturbations by leveraging the structural properties of the stationary moment equations. A key step in this development is addressing the underdetermined nature of moment equations, which has traditionally made it difficult to characterize how stationary moments depend on system parameters. To overcome this, we utilize the Dulmage-Mendelsohn (DM) decomposition of the coefficient matrix to extract welldetermined subequations and reveal their hierarchical structure. Leveraging this structure, we identify stationary moments whose partial derivatives with respect to parameters are structurally zero, facilitating the exploration of fundamental constraints that govern homeostatic behavior in stochastic biochemical systems.
@@ I Introduction
Homeostasis is the property by which organisms adapt to changes in external environments, enabling survival in highly uncertain conditions. Experimental studies reported that homeostasis also functions at the cellular scale within chemical reaction networks[1, 2, 3]. However, many underlying principles of homeostasis remain unclear, highlighting the need for model-based analysis. Mathematically, homeostasis can be described as the ability of the reaction network to maintain its state variables invariant to perturbations in parameters such as reaction rates, which may be affected by environmental factors such as temperature and osmorarity [4, 1, 5]. To date, many efforts have been made to analyze the invariance based on deterministic models, specifically the rate equations. These include approaches that utilize reaction stoichiometry to identify subnetworks attaining perfect adaptation [6, 7, 8] and those employing sensitivity analysis based on the Jacobian matrix [9, 10, 11]. However, since intracellular reactions are inherently stochastic due to the low copy number of molecules, developing methods based on stochastic models is desirable for a more accurate analysis that better aligns with experimental observations.
@@
The governing equation for stochastic chemical reactions is the chemical master equation (CME), an infinite-dimensional linear differential equation for a discrete-state Markov process that poses challenges for sensitivity and invariance analysis [12, 13]. To address this, various numerical methods were proposed to evaluate the sensitivity of the density distributions of the copy numbers using the Fisher information matrix (FIM), employing Monte Carlo simulations or model approximations [14, 15]. Meanwhile, moment-based approaches are amenable to analytic or algebraic characterization of stationary moments based on the moment equation [16, 17] and enabled rigorous analysis of invariant moments for a certain class of biomolecular networks. A notable example is the antithetic integral feedback (AIF) network, where the stationary mean copy numbers can be explicitly given in terms of biochemical parameters, allowing the theoretically guaranteed design of perfect adaptation for stochastic chemical reactions [18]. However, no existing methods achieve both (i) a mathematically rigorous analysis of parameter invariance and (ii) general applicability to arbitrary chemical reaction networks for stochastic systems. This motivates the development of a systematic approach for analyzing parameter invariance in stochastic chemical systems that can rigorously address these limitations.
@@
In this study, we propose a systematic method for analyzing the parameter invariance of stationary moments in general chemical reaction networks by leveraging the structural information of the moment equations. We utilize the Dulmage-Mendelsohn (DM) decomposition [19, 20, 21], a graph-based matrix operation that extracts a welldetermined subequations from a system of linear equations, to identify determinable variables in the highly underdetermined stationary moment equations and the dependency structure among the variables. By linking this structure to unreachability analysis in directed graphs, we establish a general algorithm to systematically identify stationary moments whose partial derivatives with respect to parameters are structurally zero. The proposed method contributes to a broader understanding of homeostasis while explicitly accounting for the stochastic nature of biochemical reactions.
@@
Notations: The symbol $\mathbb{P}(\cdot)$ is a probability mass function, and $\mathbb{E}[f(\cdot)]$ is the moment of the mass function defined by $\sum_{\bm{x}}f(\bm{x})\mathbb{P}(\bm{x})$. The expression $\bm{x}^{\bm{\alpha}}$ with $\bm{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{n}]^{\top}$ denotes multivariate monomial of $x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\dots x_{n}^{\alpha_{n}}$.
@@ II-A Chemical Master Equations
We consider chemical reaction networks $\Gamma$ consisting of $n$ molecular species $X_{i}\>(i=1,2,\dots,n)$ and $m$ reaction species $R_{j}\>(j=1,2,\dots,m)$. In the modeling of stochastic chemical reactions, the state of the reaction networks $\Gamma$ is characterized by the copy number of molecular species $X_{i}$, and its evolution is modeled by the Markov process governed by stochastic increase and decrease of the copy numbers by reactions $R_{j}$.
To formally define this stochastic process, let $x_{i}$ denote the copy number of the molecular species $X_{i}$, and define the state of the reaction network $\Gamma$ at time $t$ by $\bm{x}(t)=[x_{1},x_{2},\dots,x_{n}]^{\top}\in\mathbb{N}_{0}^{n}$. Reaction species $R_{j}$ is characterized by a stoichiometry vector $\bm{s}_{j}\in\mathbb{Z}^{n}$, which describes the change in $\bm{x}$ upon the occurrence of the reaction, and a propensity function $w_{j}(\bm{x})$ representing the probability of its occurrence per unit time. The propensity function $w_{j}(\bm{x})$ is defined for each reaction type by
$\!w_{j}(\bm{x}):=\begin{cases}\theta_{j}&\text{zero-th order reaction}\\ \theta_{j}x_{i}&\text{monomolecular reaction}\\ \theta_{j}x_{i}x_{l}&\text{bimolecular of $X_{i}$ and $X_{\ell}$}\\ {\theta_{j}x_{i}(x_{i}-1)}/2&\text{bimolecular of $X_{i}$'s}\\ \end{cases}$
with the rate parameters $\theta_{j}\in\mathbb{R}_{+}$, assuming that $R_{j}$ is one of the elementary reactions [5].
The stochastic chemical reactions are then modeled as the evolution of the copy number distribution of the state $\bm{x}$ at time $t$ denoted by $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$, for a given initial state $\bm{x}_{0}$, i.e., $\bm{x}_{0}:=\bm{x}(0)$. The governing equation of the dynamics is called the Chemical Master Equation (CME) [12, 13], or Kolmogorov’s forward equation. Omitting $t$ and $\bm{x}_{0}$ in $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$ to avoid notational clutter, the CME can be expressed as
@@ II-B Problem statement
To formally define the problem, we introduce the following assumption.
@@
This assumption guarantees the existence of the moments of the stationary copy number distribution, or the stationary moments. In what follows, we denote the stationary moments by $\mu_{\bm{\alpha}}(\bm{\theta}):=\mathbb{E}[\bm{x}^{\bm{\alpha}}|\bm{\theta}]$, where $\bm{\alpha}\in\mathbb{N}_{0}^{m}$ is a multi-index specifying the moment (see Notations in Section I). Then, the problem is formally stated as follows.
@@
For a given chemical reaction network $\Gamma$ and a reaction species $R_{k}$ to be perturbed, assume that Assumption 1 holds. Define the variation of the parameter associated with the reaction species $R_{k}$ around $\bm{\theta}$ by
@@ III-A Overview of the proposed approach
Let $\mu^{(1)}_{\bm{\alpha}}(\bm{\theta})$ denote the partial derivative of the stationary moment with respect to $\theta_{k}$ defined as
$\mu^{(1)}_{\bm{\alpha}}(\bm{\theta}):=\frac{\partial\mu_{\bm{\alpha}}}{\partial\theta_{k}}\bigg{|}_{\bm{\theta}}.$ (4)
By definition, the stationary moment $\mu_{\bm{\alpha}}(\bm{\theta})$ is invariant under the change in the parameter $\theta_{k}$, i.e., $\bm{\alpha}\in\mathcal{M}_{k}$ holds, if and only if its derivative satisfies $\mu^{(1)}_{\bm{\alpha}}(\bm{\theta})=0$ for all $\bm{\theta}$. In other words, our problem reduces to identifying the moments that remain invariant due to the structural properties of the system, independent of the specific parameter values $\bm{\theta}$.
@@
$\sum_{\bm{\beta}}\left\{a^{k}_{\bm{\alpha},\bm{\beta}}\mu_{\bm{\beta}}(\bm{\theta})+b^{k}_{\bm{\alpha}}+\sum_{j=1}^{m}\theta_{j}a^{j}_{\bm{\alpha},\bm{\beta}}\mu^{(1)}_{\bm{\beta}}(\bm{\theta})\right\}=0,$ (6)
@@
To address this challenge, we present an approach that utilizes the structural information of these algebraic equations to identify the invariant moments satisfying $\mu_{\bm{\alpha}}^{(1)}(\bm{\theta})=0$ for all $\bm{\theta}\in\mathbb{R}_{+}^{m}$. To this end, we first replace the variables (moments) in eq. (5) and eq. (6) with independent variables and define the following linear equations
$\!\!\begin{bmatrix}\sum_{j=1}^{m}\theta_{j}A_{j}&A_{k}\\ O&\sum_{j=1}^{m}\theta_{j}A_{j}\end{bmatrix}\begin{bmatrix}\bm{\nu}^{(1)}\\ \bm{\nu}\end{bmatrix}=-\begin{bmatrix}\bm{b}_{k}\\ \sum_{j=1}^{m}\theta_{j}\bm{b}_{j}\end{bmatrix},$ (7)
where $A_{j}\in\mathbb{R}^{M\times N}(j=1,2,\ldots,m)$ and $\bm{b}_{j}\in\mathbb{R}^{M}(j=1,2,\ldots,m)$ are matrices and vectors that contain the coefficients $a_{\bm{\alpha},\bm{\beta}}^{j}$ and $b_{\bm{\alpha}}^{j}$, respectively, and $\bm{\nu}$ and $\bm{\nu}^{(1)}$ are the vectors of independent variables associated with the stationary moments and their derivatives. For simplicity, we make the following assumptions that eliminate redundant equations and unnecessary variables.
@@
For all $\bm{\theta}\in\mathbb{R}^{m}_{+}$, the matrix $\sum_{j=1}^{m}\theta_{j}A_{j}$ in eq. (7) is full row rank, and at least one entry of for each column of $\sum_{j=1}^{m}\theta_{j}A_{j}$ is non-zero.
@@
The full row rank condition implies that the moment equation does not have redundant equations, which can be systematically eliminated by utilizing the basis vectors common to the left null spaces of the augmented coefficient matrices $[A_{j}\ \bm{b}_{j}]\>(j=1,2,\dots,m)$. The latter half of the assumption ensures that the vector $[\bm{\nu}^{(1)},\bm{\nu}]^{\top}$ does not contain unused variables.
Next, to identify the moments satisfying $\bm{\mu}_{\bm{\alpha}}^{(1)}(\bm{\theta})=0$ for all $\bm{\theta}$, we focus on the fact that the arrangement of non-zero elements in the matrices $A_{j}$ and the vectors $\bm{b}_{j}$ are independent of $\bm{\theta}$. In particular, we utilize the zero and non-zero structure of eq. (7) and explore the entries of $\bm{\nu}^{(1)}$ that are forced to be zero due to the structural constraint. For this purpose, we define an operator $\pi:\mathbb{R}^{M\times N}\cup\mathbb{R}^{M}\rightarrow\{0,1\}^{M\times N}\cup\{0,1\}^{M}$ that maps a matrix or vector to a binary matrix or vector encoding the zero and non-zero structure, respectively. Specifically, for a matrix $X\in\mathbb{R}^{M\times N}$ and a vector $\bm{x}\in\mathbb{R}^{M}$,
$\!\!\!(\pi(X))_{p,q}\!:=\!\!\begin{cases}\!1\!\!\!\!&\!\text{if $(X)_{p,q}\!\neq\!0$}\\ \!0\!\!\!\!&\!\text{otherwise}\end{cases}\!,(\pi(\bm{x}))_{p}\!:=\!\!\begin{cases}\!1\!\!\!&\!\!\text{if $(\bm{x})_{p}\!\neq\!0$}\\ \!0\!\!\!&\!\!\text{otherwise}\end{cases}.\!\!$
Our approach is then to seek invariant moments for all possible equations that share the same zero and non-zero structure as the stationary moment equations. The following Proposition is instrumental for the analysis of invariant moments.
Proposition 1
Consider linear equations
$\Phi\bm{\nu}_{\Phi}=\bm{b}_{\Phi},$ (8)
where $\Phi\in\mathbb{R}^{2M\times 2N}$ and $\bm{b}_{\Phi}\in\mathbb{R}^{2M}$ are the coefficient matrix and vector that share the same non-zero structure as the coefficient matrix and vector eq. (7), i.e.,
$\Phi:=\begin{bmatrix}\tilde{A}&\tilde{A}_{k}\\ O&\tilde{A}\end{bmatrix},\bm{\nu}_{\Phi}$ $:=\begin{bmatrix}\tilde{\bm{\nu}}^{(1)}\\ \tilde{\bm{\nu}}\end{bmatrix},\bm{b}_{\Phi}:=\begin{bmatrix}\tilde{\bm{b}}_{k}\\ \tilde{\bm{b}}\end{bmatrix}$ (9)
with $\tilde{\bm{\nu}}^{(1)}$ and $\tilde{\bm{\nu}}$ being vectors with the same size as ${\bm{\nu}}^{(1)}$ and ${\bm{\nu}}$, respectively, $\tilde{A}_{j}\in\{X\in\mathbb{R}^{M\times N}\mid\pi(X)=\pi({A}_{j})\},\tilde{A}\in\{X\in\mathbb{R}^{M\times N}\mid\pi(X)=\bigcup_{j=1}^{m}\pi({A}_{j}),\text{rank}(X)=M\},\tilde{\bm{b}}_{j}\in\{\bm{x}\in\mathbb{R}^{M}\mid\pi(\bm{x})=\pi(\bm{b}_{j})\}$ and $\tilde{\bm{b}}\in\{\bm{x}\in\mathbb{R}^{M}\mid\pi(\bm{x})=\bigcup_{j=1}^{m}\pi(\bm{b}_{j})\}$. The union of the binary matrices and vectors are defined so that the entry becomes 1 if at least one of the matrices and vectors has a non-zero entry in the corresponding position, and 0 otherwise, respectively. If the $i$-th entry of the solution $\bm{\nu}_{\Phi}$ is zero for every solution of eq. (8), the $i$-th entry of $[\bm{\nu}^{(1)},\bm{\nu}]^{\top}$ is also zero for all $\theta_{j}\in\mathbb{R}_{+}\ (j=1,2,\ldots,m)$ and for every solution of eq. (7).
This proposition implies that the partial derivative (4) is zero if the corresponding entry of $\bm{\nu}^{(1)}$ is zero for all possible equations having the same non-zero structure as eq. (7). In the next subsections, we show an approach that systematically identifies such entries of $\bm{\nu}^{(1)}$ of eq. (7).
III-B Dulmage-Mendelsohn decomposition
We analyze the solutions of the linear equations (8) by systematically permuting the rows and columns of the coefficient matrix $\Phi$ to make triangular blocks corresponding to welldetermined and underdetermined subequations. The entries of the solution $\bm{\nu}_{\Phi}$ that are structurally constrained to be zero are then obtained from the welldetermined subequations. For this purpose, we utilize the Dulmage-Mendelsohn (DM) decomposition [19, 20, 21], originally introduced as a partitioning algorithm for bipartite graphs. The following lemma shows the existence and the uniqueness of such block-triangular decomposition of the matrix $\tilde{A}$, which can be systematically implemented as a maximum matching problem of a bipartite graph[19].
Lemma 1
[20] Consider the matrix $\tilde{A}$ in eq. (9). There exist row and column permutation matrices $P$ and $Q$ that transform $\tilde{A}$ into the upper-triangular block matrix by
${P}\tilde{A}{Q}=\begin{bmatrix}\tilde{A}_{u}&\tilde{A}_{ud}\\ O&\tilde{A}_{d}\end{bmatrix},$ (10)
where $\tilde{A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block-triangular matrix with $\eta$ diagonal blocks satisfying $\text{rank}(\tilde{A}_{d})=N_{d}$, and $\tilde{A}_{u}\in\mathbb{R}^{(M-N_{d})\times(N-N_{d})}$ is a matrix satisfying $\text{rank}(\tilde{A}_{u})=M-N_{d}$. Among such transformations, the partition maximizing $\eta$ is unique, and each diagonal block in $\tilde{A}_{d}$ is uniquely determined up to row and column permutations within that block.
It is known that the DM decomposition of eq. (10) produces such block-triangular form that maximally separates the welldetermined component with the finest partitions from the original matrix [20]. This allows for extracting the moments and their derivatives that can be fully determined from the underdetermined moment equations. The next lemma shows that the maximal welldetermined block of $\Phi$ in eq. (9) is obtained using $\tilde{A}_{d}$ in Lemma 1.
Lemma 2
There exist row and column permutations that transform $\Phi$ in eq. (9) into the upper-triangular block matrix
$\Psi=\begin{bmatrix}\tilde{A}_{u1}&\tilde{A}_{u2}&\tilde{A}_{u3}\\ O&\tilde{A}_{d}&\tilde{A}^{k}_{d}\\ O&O&\tilde{A}_{d}\end{bmatrix},$ (11)
where $\tilde{A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block-triangular matrix with $\eta$ diagonal blocks defined in eq. (10), and $\tilde{A}_{d}^{k}\in\mathbb{R}^{N_{d}\times N_{d}},\tilde{A}_{u1}\in\mathbb{R}^{2(M-N_{d})\times 2(N-N_{d})},\tilde{A}_{u2}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ and $\tilde{A}_{u3}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ are real-valued matrices defined accordingly. Moreover, this decomposition maximizes the diagonal blocks among all row and column permutations of $\Phi$, and each diagonal block of $\Psi$ is unique up to row and column permutations within the block.
The proof of this proposition is shown in Appendix -A. Using Lemma 2, we rearrange eq. (8) to obtain
$\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi},$ (12)
where $\bm{\nu}_{\Psi}\in\mathbb{R}^{2N}$ and $\bm{b}_{\Psi}\in\mathbb{R}^{2M}$ are the vectors obtained by rearranging the entries of $\bm{\nu}_{\Phi}$ and $\bm{b}_{\Phi}$ according to the transformation from the matrix $\Phi$ to the matrix $\Psi$. Consequently, all welldetermined equations are extracted with the block-triangular coefficient matrix $\Psi$ as illustrated in Fig. 1, where $\bullet$ represents an arbitrary non-zero element. Thus, by applying backward substitution for the linear equations (12), we can systematically determine the entries of $\bm{\nu}_{\Psi}$ that are zero regardless of the coefficients, from which invariant moments are identified.
Figure 1: Illustration of a matrix transformation that extracts a welldetermined upper-triangular equation (shown in green). In this example, $\mathbb{E}[x_{3}]^{(1)}=0$ holds from the structural information of the equation.
@@ III-B Dulmage-Mendelsohn decomposition
Example: Consider the reaction network with molecular sequestrations $\Gamma$, where the stoichiometry vector $\bm{s}_{j}$ and the propensity function $w_{j}(\bm{x})$ are given in Table I.

### v2

@@ Abstract
Living organisms maintain stable functioning amid environmental fluctuations through homeostasis, a property that preserves a system’s behavior despite changes in environmental conditions. To elucidate homeostasis in stochastic biochemical reactions, theoretical tools for assessing population-level invariance under parameter perturbations are crucial. In this paper, we propose a systematic method for identifying the stationary moments that remain invariant under parameter perturbations by leveraging the structural properties of the stationary moment equations. A key step in this development is addressing the underdetermined nature of moment equations, which has traditionally made it difficult to characterize how stationary moments depend on system parameters. To overcome this, we utilize the Dulmage-Mendelsohn (DM) decomposition of the coefficient matrix to extract welldetermined subequations and reveal their hierarchical structure. Leveraging this structure, we identify stationary moments whose partial derivatives with respect to parameters are structurally zero, facilitating the exploration of fundamental constraints that govern homeostatic behavior in stochastic biochemical systems.
@@ I Introduction
Homeostasis is the property by which organisms adapt to changes in external environments, enabling survival in highly uncertain conditions. Experimental studies have begun to reveal the underlying structural mechanisms of biomolecular reactions that enable homeostatic behaviors [1, 2, 3]. In this line of research, a recent study developed a synthetic biomolecular feedback system that is robust to disturbances [4] based on the theoretical analysis of stochastic biomolecular reaction networks. Mathematically, homeostasis can be described as the ability of the reaction network to maintain its state variables invariant to perturbations in parameters such as reaction rates, which may be affected by environmental factors such as temperature and osmolarity [4, 1, 5]. To date, many efforts have been made to analyze the invariance based on deterministic models, specifically the rate equations. These include approaches that utilize reaction stoichiometry to identify subnetworks attaining perfect adaptation[6, 7] and those employing sensitivity analysis based on the Jacobian matrix [8, 9]. However, since intracellular reactions are inherently stochastic due to the low copy number of molecules, developing methods based on stochastic models is desirable for a more accurate analysis that better aligns with experimental observations.
@@
The governing equation for stochastic chemical reactions is the chemical master equation (CME), an infinite-dimensional linear differential equation for a discrete-state Markov process that poses challenges for sensitivity and invariance analysis [10, 11]. To address this, various numerical methods were proposed to evaluate the sensitivity of the density distributions of the copy numbers using the Fisher information matrix (FIM), employing Monte Carlo simulations or model approximations [12, 13]. Meanwhile, moment-based approaches are amenable to analytic or algebraic characterization of stationary moments based on the moment equation [14, 15, 5] and enabled rigorous analysis of invariant moments for a certain class of biomolecular networks. A notable advancement in this direction is the mathematical characterization of the antithetic integral feedback (AIF) network [16] as a system that achieves maximal robust perfect adaptation (MaxRPA), a property in which the stationary moments are robust to all but two parameter perturbations [17]. Yet, a general framework is still lacking for systematically identifying stationary moments that remain invariant under arbitrary parameter perturbations. Addressing this gap is essential for advancing the theoretical understanding of homeostasis beyond known motifs.
@@
In this study, we propose a systematic method for analyzing the invariance of stationary moments with respect to arbitrary parameters by leveraging the structural information of the moment equations. We utilize the Dulmage-Mendelsohn (DM) decomposition [18, 19], a graph-based matrix operation that extracts welldetermined subequations from a system of linear equations, to identify determinable variables in the highly underdetermined stationary moment equations and the dependency structure among the variables. By linking this structure to unreachability analysis in directed graphs, we establish a computationally tractable algorithm to systematically identify stationary moments whose partial derivatives with respect to parameters are structurally zero. The proposed method contributes to a broader understanding and application of homeostasis by explicitly accounting for the stochastic nature of biochemical reactions.
@@
Notations: The symbol $\mathbb{P}(\cdot)$ is a probability mass function, and $\mathbb{E}[f(\cdot)]$ is the moment of the mass function defined by $\sum_{\bm{x}}f(\bm{x})\mathbb{P}(\bm{x})$. The symbol $\mathrm{supp}(\cdot)$ denotes the set of non-zero entries, i.e., $\{(i,j)\mid(X)_{i,j}\neq 0\}$ for a matrix $X$ and $\{i\mid(\bm{x})_{i}\neq 0\}$ for a vector $\bm{x}$, respectively. The expression $\bm{x}^{\bm{\alpha}}$ with $\bm{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{n}]^{\top}$ denotes a multivariate monomial of $x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\dots x_{n}^{\alpha_{n}}$.
@@ II-A Chemical Master Equations
We consider chemical reaction networks $\Gamma$ consisting of $n$ molecular species $X_{i}\>(i=1,2,\dots,n)$ and $m$ reaction species $R_{j}\>(j=1,2,\dots,m)$. In stochastic chemical reactions, the state of the reaction networks $\Gamma$ is characterized by the copy number of molecular species $X_{i}$, and its evolution is modeled by the Markov process governed by stochastic changes in the copy numbers by reactions $R_{j}$.
To formally define this stochastic process, let $x_{i}$ denote the copy number of the molecular species $X_{i}$, and define the state of the reaction network $\Gamma$ at time $t$ by $\bm{x}(t)=[x_{1},x_{2},\ldots,x_{n}]^{\top}\in\mathbb{N}_{0}^{n}$. Reaction species $R_{j}$ is characterized by a stoichiometry vector $\bm{s}_{j}\in\mathbb{Z}^{n}$, which describes the change in $\bm{x}$ upon the occurrence of the reaction, and a propensity function $w_{j}(\bm{x})$ representing the probability of its occurrence per unit time. The reaction $R_{j}$ is one of the elementary reactions, and the propensity function $w_{j}(\bm{x})$ is linear with respect to $\theta_{j}$ and a polynomial in $\bm{x}$ with a maximum degree of two [5].
The stochastic chemical reactions are then modeled as the evolution of the copy number distribution of the state $\bm{x}$ at time $t$ denoted by $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$, for a given initial state $\bm{x}_{0}$, i.e., $\bm{x}_{0}:=\bm{x}(0)$. The governing equation of the dynamics is called the Chemical Master Equation (CME) [10, 11]. Omitting $t$ and $\bm{x}_{0}$ in $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$ to avoid notational clutter, the CME can be expressed as
@@ II-B Problem statement
To formally define the problem, we introduce the following assumption, following literature [15].
@@
This assumption ensures the existence of a stationary distribution and stationary moments of molecular copy numbers, which is closely related to the ergodicity of the underlying Markov process [20]. The ergodicity of stochastic chemical reaction networks can be verified using tractable optimization algorithms [21]. In what follows, we denote the stationary moments by $\mu_{\bm{\alpha}}(\bm{\theta}):=\mathbb{E}[\bm{x}^{\bm{\alpha}}|\bm{\theta}]$, where $\bm{\alpha}\in\mathbb{N}_{0}^{m}$ is a multi-index specifying the moment (see Notations in Section I). Then, the problem is formally stated as follows.
@@
For a given chemical reaction network $\Gamma$ and one of the reaction species $R_{k}$, assume that Assumption 1 holds. Define the variation of the parameter associated with the reaction species $R_{k}$ around $\bm{\theta}$ by
@@ III-A Overview of the proposed approach
Let $\partial_{k}\mu_{\bm{\alpha}}(\bm{\theta})$ denote the partial derivative of the stationary moment with respect to $\theta_{k}$ defined as
$\partial_{k}\mu_{\bm{\alpha}}(\bm{\theta}):=\frac{\partial\mu_{\bm{\alpha}}}{\partial\theta_{k}}\bigg{|}_{\bm{\theta}}.$ (4)
By definition, the stationary moment $\mu_{\bm{\alpha}}(\bm{\theta})$ is invariant under the change in the parameter $\theta_{k}$, i.e., $\bm{\alpha}\in\mathcal{M}_{k}$ holds, if and only if $\partial_{k}\mu_{\bm{\alpha}}(\bm{\theta})=0$ for all $\bm{\theta}$. In other words, our problem reduces to identifying the moments that remain invariant due to the structural properties of the system, independent of the specific parameter values $\bm{\theta}$.
@@
$\sum_{\bm{\beta}}\left\{a^{k}_{\bm{\alpha},\bm{\beta}}\mu_{\bm{\beta}}(\bm{\theta})+b^{k}_{\bm{\alpha}}+\sum_{j=1}^{m}\theta_{j}a^{j}_{\bm{\alpha},\bm{\beta}}\partial_{k}\mu_{\bm{\beta}}(\bm{\theta})\right\}=0,$ (6)
@@
To address this challenge, we present an approach that utilizes the structural information of these algebraic equations to identify the invariant moments satisfying $\partial_{k}\mu_{\bm{\alpha}}(\bm{\theta})=0$ for all $\bm{\theta}\in\mathbb{R}_{+}^{m}$. To this end, we first replace the variables (moments) in eq. (5) and eq. (6) with independent variables and define the following linear equations
$\Phi(\bm{\theta})\bm{\nu}_{\Phi}$ $=\bm{b}_{\Phi}(\bm{\theta}),$ (7)
where
$\!\!\!\Phi(\bm{\theta})\!:=\!\begin{bmatrix}A(\bm{\theta})\!\!\!&\!\!\!A_{k}\\ O\!\!\!&\!\!\!A(\bm{\theta})\end{bmatrix},\bm{\nu}_{\Phi}\!:=\!\begin{bmatrix}\partial_{k}\bm{\nu}\\ \bm{\nu}\end{bmatrix},\bm{b}_{\Phi}(\bm{\theta})\!:=\!-\begin{bmatrix}\bm{b}_{k}\\ \bm{b}(\bm{\theta})\end{bmatrix}\!$ (8)
with $A(\bm{\theta}):=\sum_{j=1}^{m}\theta_{j}A_{j}$, $\bm{b}(\bm{\theta}):=\sum_{j=1}^{m}\theta_{j}\bm{b}_{j}$, $A_{j}\in\mathbb{R}^{M\times N}$ and $\bm{b}_{j}\in\mathbb{R}^{M}\>(j=1,2,\dots,m)$ are matrices and vectors containing the coefficients $a_{\bm{\alpha},\bm{\beta}}^{j}$ and $b_{\bm{\alpha}}^{j}$, respectively, and the vectors $\bm{\nu}$ and $\partial_{k}\bm{\nu}$ denote independent variables associated with the stationary moments and their derivatives. For simplicity, we make the following assumptions that eliminate redundant equations and unnecessary variables.
@@
For all $\bm{\theta}\in\mathbb{R}^{m}_{+}$, the matrix $A(\bm{\theta})$ in eq. (7) is full row rank, and at least one entry of each column of $A(\bm{\theta})$ is non-zero.
@@
The full row rank condition implies that the moment equation does not have redundant equations, which can be systematically eliminated by utilizing the basis vectors common to the left null spaces of the augmented coefficient matrices $[A_{j}\ \bm{b}_{j}]\>(j=1,2,\dots,m)$. The latter half of the assumption ensures that the vectors $\bm{\nu}_{\Phi}$ do not contain unused variables.
In the rest of this section, we illustrate the overview of the proposed approach using a simple example.
@@
Example: Consider the reaction network with molecular sequestration $\Gamma$, where the stoichiometric vectors $\bm{s}_{j}$ and propensity functions $w_{j}(\bm{x})$ are defined in Table I.